### v1

# EvAlignUX: Advancing UX Research through LLM-Supported Exploration of Evaluation Metrics

Abstract.
Evaluating UX in the context of AI’s complexity, unpredictability, and generative nature presents unique challenges. HCI scholars lack sufficient tool support to build knowledge around diverse evaluation metrics and develop comprehensive UX evaluation plans. In this paper, we introduce EvAlignUX , an innovative system grounded in scientific literature and powered by large language models (LLMs), designed to help HCI scholars explore evaluation metrics and their relationship to potential research outcomes. A user study involving 19 HCI scholars revealed that EvAlignUX significantly improved the perceived clarity, specificity, feasibility, and overall quality of their evaluation proposals. The use of EvAlignUX enhanced participants’ thought processes, resulting in the creation of a Question Bank that can be used to guide UX Evaluation Development. Additionally, the influence of researchers’ backgrounds on their perceived inspiration and concerns about over-reliance on AI highlights future research directions for AI’s role in fostering critical thinking.
1. Introduction
The rapid integration of AI into interactive systems presents a growing challenge for user experience (UX) evaluation (Stige et al., 2023; Chaudhry, 2024). Traditional usability evaluation guided by standard frameworks such as the System Usability Scale (SUS) (Bangor et al., 2008), User Experience Questionnaire (UEQ) (Schrepp et al., 2017), NASA Task Load Index (NASA-TLX) (Hart, 2006), fall short in capturing the complex, dynamic interaction between humans and AI (Kuang et al., 2024). As AI systems become more context-aware and capable of generating responses in real-time, UX researchers are confronted with new questions about trust, emotional engagement, and ethical considerations (Liao et al., 2020a; Ahmad et al., 2023a; Shneiderman, 2020; Ntoa et al., 2021). Addressing these new considerations is especially challenging for systems research, where a range of evaluation methods, often borrowing and combining techniques from software engineering, design, and usability evaluation, is triangulated to evaluate the UX of interactive computing systems (Marquardt et al., 2017). Current UX evaluation methods, e.g., Post-Study System Usability Questionnaire (PSSUQ) (Lewis, 1992), Subjective Mental Effort Questionnaire (SMEQ) (Hyland et al., 1989), while robust for traditional systems, are inadequate for AI-powered interfaces. Researchers are striving to develop new frameworks, which take into account ethical considerations, trust dynamics, intimacy, and multimodal behaviors, to adequately guide scholars and practitioners in developing their UX evaluation plans for human-AI interaction research, e.g., (Zheng et al., 2022; Liao et al., 2024). Given the plethora of well documented evaluation methods, heuristics and metrics (e.g., (Lazar et al., 2017; Nielsen, 1992; Vermeeren et al., 2010)), much less is known about How researchers select appropriate metrics to conduct a comprehensive evaluation of a system.
Also, despite the availability of various UX evaluation methods (Pettersson et al., 2018; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019; Kieffer et al., 2019), there is a lack of practical tools to guide researchers in developing comprehensive evaluation plans. Research has shown that incorporating evaluation metrics as part of the design materials in early stages can improve the clarity of designing human-AI interaction research (Zheng and Huang, 2023). UX is frequently regarded as a rigid set of methods, leading to calls for research and education that encourage framing UX as “more of a mindset than a method” within the community (Gray, 2016). However, with the rapid advancement of AI models, the literature on AI-mediated multi-stakeholder interactions has expanded exponentially. This growth presents challenges for novice researchers in building up knowledge of relevant evaluation metrics, while seasoned researchers also face increasing pressure to stay current with the latest developments.
To bridge this gap, we designed, implemented, and evaluated EvAlignUX , an interactive system that supports researchers in developing UX evaluation plans for their research projects. EvAlignUX is powered by the generative capabilities of Large Language Models (LLMs) to facilitate the exploration of UX evaluation ideas, as LLMs have demonstrated their potential for ideation of scientific research by processing a vast amount of information, generating context-aware insights, retrieving literature, and brainstorming (Doshi and Hauser, 2024; Glickman and Zhang, 2024; Liu et al., 2024). Specifically, as shown in Fig. 1, EvAlignUX comprises three primary panels: the Project Ideation Panel, the Metric Explorer Panel, and the Outcomes and Risks Panel. These panels allow users to explore UX metrics, potential outcomes, and potential risks associated with human-AI interaction research. In an evaluation with 19 HCI scholars, EvAlignUX led to significant improvements in the clarity, concreteness, and confidence in their UX proposals. The results demonstrated that EvAlignUX significantly improved the clarity, concreteness, and confidence of researchers in their proposed UX evaluation plans. Participants were able to identify and integrate detailed UX metrics, self-critique their plans, and raise awareness of potential risks using the system features.
This work makes several significant contributions to the HCI community. First, EvAlignUX is a novel system, designed to support the ideation process of developing UX evaluation plans. Powered by LLMs, EvAlignUX enables users to explore UX metrics and their relationships with potential research outcomes. The unique visual interface of the metrics graph supports users in examining the connections between different UX metrics. To the best of our knowledge, EvAlignUX fills the gap in tool support for researchers, providing insights into the research literature of human-AI evaluation metrics and facilitating better decision-making regarding UX evaluation. Second, our user studies demonstrate the effectiveness of EvAlignUX in helping users gain significantly more clarity and feasibility in their proposed UX evaluation plans. Our findings revealed that researchers’ background impacted their perceived levels of inspiration and their concerns about over-reliance on AI. They also integrate the system features into their thought processes differently, focusing on varied utilities such as comprehensiveness of self-critique and efficiency of implementing the ideas. Third, our analysis of participants’ think-aloud and perceived value of different features produced a comprehensive Question Bank, highlighting the specific questions researchers asked at different steps of UX evaluation development and identifying how AI provide valuable support such as envisioning the research impact, assessing the risks involved in the research, and justifying metric selection. Lastly, our findings prompt a timely and profound discussion on shifting the focus from method to mindset in future research on UX evaluation.
2. Related Work
In this section, we review the literature on (1) current UX evaluation approaches, and (2) the potential of leveraging LLMs to provide tool support.
2.1. UX Evaluation of HCI System Research
Developing a UX evaluation plan can be challenging, as it’s often unclear what constitutes ”evaluation” and which metrics are appropriate (Ledo et al., 2018; Hillman et al., 2023). Hillman et al. (Hillman et al., 2023) elaborated on several challenges involved in developing a solid evaluation plan, such as selecting metrics that accurately reflect UX and research goals, choosing data collection methods, establishing collection protocols, and ensuring consistency across products and organizations. Additionally, long-term effects and social or cultural barriers further complicate the process (Hillman et al., 2023; Darin et al., 2019; Vermeeren et al., 2010). In the past, HCI scholars often relied on established surveys like System Usability Scale (SUS), User Experience Questionnaire (UEQ), Post-Study System Usability Questionnaire (PSSUQ), NASA Task Load Index (NASA-TLX), Subjective Mental Effort Questionnaire (SMEQ) and Customer Satisfaction (CSAT) to measure users’ perceptions (Bangor et al., 2008; Schrepp et al., 2017; Lewis, 1992; Hart, 2006; Reichheld, 2003). PULSE metrics (Page views, Uptime, Latency, etc.) are also commonly used to track user behavior (Rodden et al., 2010). However, these metrics are often “low-level or indirect measures of user experience” (Rodden et al., 2010). Several evaluation frameworks relating to user-centered metrics have been introduced to the HCI and UX domains, such as the Software Development Productivity Framework (Sadowski et al., 2019), and Google’s HEART framework (Rodden et al., 2010). For instance, the HEART framework, which includes Happiness, Engagement, Adoption, Retention, and Task success, was introduced for measuring web-based applications at scale (Rodden et al., 2010).
With the rise of AI, recent studies on UX research lay out several new trends and challenges. A school of HCI researchers aims to identify and propose new evaluation metrics for both usability and user-centered evaluation in the age of AI. For example, Zheng et al.’s literature review of human-AI interaction research in HCI, which included user evaluations, reported a suite of metrics using a survey, system logs, and interviews to collect (Zheng et al., 2022) and argued the importance of including more user-centered metrics (e.g., sociability, socio-emotional, and AI anthropomorphism metrics) with AI’s evolving role in human social interactions. Likewise, Faruk et al.’s works also proposed new subjective scales for measuring the UX of voice agents (Faruk et al., 2024a) and conversational agents (Faruk et al., 2024b). In other domains, like healthcare, Abbasian et al. also summarized some foundation metrics for evaluating the effectiveness of healthcare conversations powered by generative AI (Abbasian et al., 2024).
However, there is a lack of tools to help researchers understand the evaluation metrics literature and design a comprehensive evaluation plan for human-AI interaction research. As Yang et al. (Yang et al., 2020) systematically reported how AI system’s uncertainty and complexity can make unpredictable errors, which easily damage UX and lead to undesired societal impact. Gray’s finding of ”More of a Mindset Than a Method” also highlighted that UX practitioners often follow a codified method rather than adopting a ”mindset” that aligns with their design process, practice context, and the specific design problem (Gray, 2016). All these works suggest that new tools and approaches are needed to support researchers and practitioners in developing their UX evaluation plans.
2.2. The Potential of LLMs to Support Ideation
Large Language Models (LLMs) can process vast amounts of multi-modal information and generate context-aware, coherent text, making them ideal tools for evaluation-related tasks in various domains. A group of researchers developed novel interaction prototypes using LLMs for evaluating generative outputs (Desmond et al., 2024; Kim et al., 2024; Shankar et al., 2024). For example, Kim et al. developed EvalLM (Kim et al., 2024), an interactive system for refining prompts by evaluating multiple outputs on user-defined criteria. This helps developers compare their prompt outputs to diagnose weaknesses in their prompts. Regarding LLMs’ support for scientific tasks, recent studies have showcased LLM-powered methods for retrieving and recommending relevant literature (Shen et al., 2023), reviewing and critiquing research drafts, writing research manuscripts, and augmenting research ideation and brainstorming for scientific writing based on scientific evidence and database (Nigam et al., 2024), utilizing large scientific dataset. Recent advancements in HCI and AI research also attempt to support UX practitioners with AI-enabled tools to conduct usability testing (Fan et al., 2022; Kuang et al., 2024), e.g., LLMs are employed to help human evaluators to identify more usability issues in usability test video recordings. All these research highlights the significant potential of LLMs to support research ideation and provide valuable feedback for task evaluation. However, little is known about how researchers with varying levels of experiences can benefit from using LLM-supported tools to develop their UX evaluation plans.
3. EvAlignUX
To address the research gap, EvAlignUX was designed as a tool to aid researchers explore evaluation metrics for developing UX plans of Human-AI research. The system was developed through an iterative process, informed by both the literature review (Zheng et al., 2022; Inan Nur et al., 2021; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019) and input from a focus group. Leveraging the authors’ expertise in HCI research and practices, we began with an initial process of developing UX evaluation plan, e.g., proposing of a research topic, identifying relevant UX metrics for evaluation, and then framing the expected research contribution. The design was refined over two months through multiple weekly feedback sessions with an HCI research group, consisting of nine graduate students and faculty members from the HCI, Informatics, and Computer Science departments (n=9). All of the participants have worked on at least one HCI project where they designed or developed systems. Their applications covered diverse contexts: vision-based question-answering, ideation and writing, accessibility, immersive technologies, and intelligent assistants. Based on the UX evaluation challenges reported from prior literature (Ledo et al., 2018; Hillman et al., 2023), and after piloting the initial pipeline with this group, we established three design goals (DG), then we iteratively searched for prior literature and consulted for feedback through this weekly focus group meetings.
DG1: Recommend UX metrics relevant for a project;
DG2: Explain how the metrics are used in existing literature; and
DG3: Inform researchers with potential risks associated with their design and evaluation plans.
3.1. System Design
Based on the design goals, EvAlignUX , finalizes with three main panels, as shown in Figure 1. To illustrate interactions in EvAlignUX , we present a walk-through with Sarah, a UX researcher designing an AI counseling chatbot for alcohol addiction. The user journey is likened to a ”shopping” process, where Sarah selects ideal evaluation plans for her proposed system.
User Journey:
Once landing on EvAlignUX , Sarah inputs her project details. She names the project “Mindful Drinking”, selects “brainstorming” and “designing the study” as the current project statuses, and provides a project description, evaluation plans,
and details the expected outcomes
After entering all details, she clicks “Generate” and proceeds to the main interaction page.
3.1.1. Project Ideation Panel (DG1)
This panel extracts relevant indexes to categorize the work within the existing literature. In the rapidly evolving field of UX research, evaluators often face challenges in understanding the context and relevance of newly proposed systems (Ledo et al., 2018). To enhance the UX evaluator’s understanding of the context in which a system operates and its significance within the broader field of UX research, we built on categorizations from prior literature reviews and key guiding literature for UX of AI design practices (Zheng et al., 2022; Liao et al., 2020b; Vermeeren et al., 2010; Kieffer et al., 2019). We leverage these established frameworks to integrate critical factors into the system output. This output is designed to automatically map the proposed system, based on the provided narrative, into categories including paradigms, application domains, modalities, system features, design novelty, design methods, human-AI relationship types, stakeholders, social scales, and theoretical frameworks.
User Journey:
Sarah views a list of indexed recommendations from EvAlignUX , noting that some are selected while others are not. She realizes ”Addiction Recovery” isn’t checked under ”Application Domain,” so she selects it (A). She also checks ”Family Members” as the evaluation target stakeholder, which she had previously overlooked. After making these changes, she clicks ”Regenerate” (B) followed by ”Compare Metrics” (C). This process updates her a list, showing added, retained, and removed metrics (D), sparking her curiosity about the new metrics.
3.1.2. Metric Explorer Panel (DG1, DG2)
This panel recommends relevant evaluation metrics and related literature while enabling users to explore additional metrics through both list-view and graph-view interactions, helping them formulate a comprehensive evaluation plan. Prior literature has highlighted that users are often excluded from the process of defining evaluation criteria for a system, with most participants consulted only at the end of the design process to assess the system’s effects, rather than being actively involved in setting these criteria. This exclusion contradicts the goal of democratizing the design process (Bossen et al., 2016; Zheng and Huang, 2023). Inspired by this challenge, we propose a metric recommendation feature that introduces UX evaluation metrics into design materials as a design or evaluation input. Given that graph-based interaction is a commonly used and effective approach for recommendation systems (Ali et al., 2020), we propose incorporating a graph-view alongside the traditional list-view. This graph-view will allow users to visualize UX metrics and their correlating metrics, fostering inspiration and enabling more informed filtering and selection of metrics. Furthermore, prior literature outlines various methods for collecting metrics, such as system logs, surveys, and interviews (Zheng et al., 2022). To enhance user understanding, we will provide context and examples of how these metrics have been derived and applied.
User Journey:
Sarah starts by reviewing suggested metrics in the list (A) and is intrigued by ”goal attainment,” defined as assessing users’ perceptions of achieving their recovery goals through the system (D). She clicks the metric info button (C) to explore how this metric is measured—through surveys, system logs (E), and interviews—and reviews related papers on each method’s application and findings (F). Interested, she adds it to her shopping cart (B) and uses the Metric Graph feature to explore additional metrics.
3.1.3. Outcomes and Risks Panel (DG2, DG3)
Assists users in framing their research contributions based on previous studies and identifies potential pitfalls by referencing databases of prior AI risks and incidents. Technologies have positively transformed society, but they have also led to undesirable consequences not anticipated at the time of design or development. Previous HCI research explored using insights into past undesirable consequences in the world to help researchers and practitioners gain awareness anticipate potential adverse effects and reflect on their own experiences with technology (Pang et al., 2024). Inspired by this, we propose a panel for generating related research outcomes and risks based on existing research and real-world cases. We sourced case studies from a database indexing real AI incidents documented by journalists — the AI, Algorithmic, and Automation Incident and Controversy (AIAAIC) repository that previous papers have used (Lee et al., 2024b). This approach aims to strengthen UX evaluation metrics by incorporating research outcomes based on literature, and relevant risks grounded in real-world incidents.
As Sarah adds and removes metrics from her shopping cart, she notices that the results in the Outcome feature update dynamically. She selects outcomes from prior studies that catch her interest. Then, she explores potential risks associated with her proposed UX evaluation plan by clicking on the Risk feature. “All set!”, said Sarah. She then proceeds with exporting the data and conducts a final review of her generated UX evaluation plan and UX outcomes report.
3.2. System Backend and Implementation
In this section, we describe the process of preparing the data repository, system backend and implemenation, as well as the techniques used for prompting and making recommendations. The frontend of the system is implemented using ReactJS111https://github.com/facebook/react, TypeScript and CSS. The backend uses Python with FastAPI222https://github.com/tiangolo/fastapi/ as the RESTful API server framework. The data is represented as a Knowledge Graph using Neo4j333https://github.com/neo4j Graph Database to model the complex relations among research papers (Verma et al., 2023; Kejriwal, 2022; Wang et al., 2018). We use the GPT-4o model as our LLM engine and set the temperature to 0.7 for all components. The embeddings in the system are generated using text-embedding-3-small, with FAISS444https://github.com/facebookresearch/faiss as the vector database for data retrieval. Additionally, Langchain555https://github.com/langchain-ai/langchain is used for prompting the GPT-4o model.
3.2.1. Data Preparation
In this section, we present how the scientific database was prepared. Based on several UX evaluation survey papers, e.g.,(Zheng et al., 2022; Inan Nur et al., 2021; Vermeeren et al., 2010), we established an initial repository of evaluation metrics. Two authors expanded this repository by incorporating additional metrics identified in the reviewed papers. Definitions for all metrics, including the newly added ones, were either directly sourced from these surveys or developed collaboratively by the authors to ensure completeness and coherence.
After building the repository of evaluation metrics, we created a paper database from the same survey papers and utilized GPT-4o (OpenAI et al., 2024) to build up more fine-grained information associated with the evaluation metrics. We prompted GPT with each full text of the papers and asked GPT to provide a narrative of the system developed in the papers, along with the challenges or the research gaps that the papers addressed. We provided index definitions and asked GPT to generate a list of ten descriptive indices based on the content of the papers. These descriptive indices include: interaction paradigms, application domain, modality, system features, design novelty, design methods, human-AI relationship types, stakeholders involved, social scale and theoretical frameworks used. The definitions for these indices can be found in Appendix A.1.
Between the metrics and the papers in the database, we also created the usage information of each metric in relevant paper. For example, a metric usage follows the following format: ”This paper uses the [X] metric to evaluate users’ [Y] towards [Z technology]. It was found that [S findings related to Y].” [X] is the metric identified in 1, [Y] is the aspect of user interaction being evaluated, [Z] is the technology or system being evaluated, and [S] is the findings related to the user interaction being evaluated.
We then crawled metadata and appended metadata to each paper. The metadata includes the uid of the paper, the title, authors, the abstract, the date of publication, uids of papers that cited this paper, keywords, the publication venue, the publisher and the affiliation of authors. Besides, we also leverage the existing AI Incident Database as a repository for risks (McGregor, 2020) and extracted relevant incidents and risks. EvAlignUX ’s backend recommends UX metrics based on an initial System Description. After a series of user interactions, it generates a new Evaluation Plan and UX Outcome using LLM and data from research papers via a Knowledge Graph, representing complex entity relations derived from source documents.
3.2.2. Index Generation and derivation of Knowledge Graph
A list of indexes is generated by EvAlignUX using the initial system description provided by the user. The indexes are intended to standardize the system description into fixed categories. Indexes are generated by providing the System Description, examples of other generated indexes, and corresponding definitions of each index as input to the LLM. Subsequently, the user is recommended three additional index values corresponding to each category. These indexes and the initial system description are used to retrieve a list of metrics from our Knowledge Graph.
The Knowledge Graph serves as the foundation for recommending metrics, built using the Neo4j Graph Database. We structured our data as a graph to capture the intricate relationships within our dataset (Verma et al., 2023; Kejriwal, 2022; Wang et al., 2018), enabling efficient information extraction and retrieval of global contextual information (Edge et al., 2024). The backing dataset of research papers is organized into distinct clusters of communities (Fortunato, 2010) based on their index similarity, citations, and common metrics using the Louvain Community detection algorithm (De Meo et al., 2011). The graph consists of three different types of nodes, each characterized by distinct properties. Table 5 in the Appendix provides a detailed overview of the distinct nodes and their associated properties. The three unique nodes are as follows:
Paper Node: This node contains details about each unique research paper.
Metric Node: The node includes attributes specific to a UX metric. This node is attached to the Paper and the Outcome node via several edges. The node is shared between two paper nodes if they measure the metric.
Outcome Node: This node is attached to the Metric node via an edge and contains the outcome and other information corresponding to each metric used in a paper. Since the metric node could be shared between multiple paper nodes, this node stores the unique information of a metric corresponding to a research paper.
All nodes are unique and do not contain duplication. Furthermore, weighted edges exist between the nodes. For instance, there exist three edges amongst each pair of paper nodes, namely,
[SHARED_METRIC], [CITES], and [CITED_BY]. The
[SHARED_METRIC] edge exists between two paper nodes if they share at least one common metric. Similarly, if any paper A cites another paper B, there exists a [CITES] edge from paper node A to paper node B, and a [CITED_BY] edge from paper node B to paper node A. Edges of type [CITES] and [CITED_BY] are assigned relatively higher weight, compared to [SHARED_METRIC] edge, by multiplying the edge weight with a constant value, due to stronger relation amongst the papers because of the citation relation, based on an increase in the modularity score as discussed in (Blondel et al., 2008). The weight assigned to these edges is influenced by the relative importance of each index category in evaluating the similarity between research papers. Index categories that indicate greater similarity receive higher weights, which are then multiplied by the similarity score between index values of the categories for each paper, contributing to the overall edge weight. Additionally, each Paper node is connected to a Metric node via a non-weighted edge of type [HAS_METRIC], which is then connected to an Outcome node via another non-weighted edge of type [HAS_OUTCOME].
3.2.3. Metrics Recommendation
The provided system description and its corresponding generated indexes extract metrics from the Knowledge Graph. The system description and indexes are vectorized, and the most similar paper is determined using cosine similarity in the existing Knowledge Graph. The community for the most similar paper is determined from the above Knowledge Graph and consequently, all metrics linked to the paper nodes belonging to the above community are extracted. These metrics, along with the user system description and generated indexes are used to further filter metrics, by prompting a LLM. Additionally, each metric is further classified into three metric collection methods, based on their usage in previous research papers, using the Knowledge Graph. These metrics are displayed in a list view and a graphical view to the user.
3.2.4. Potential risks and metric outcomes
Each metric in the Knowledge Graph is associated with outcomes corresponding to the research paper. Subsequently, these results are retrieved from the Knowledge Graph based on the final metrics list, and displayed to the user to filter, based on their relevance to the user’s system description. Each outcome is specific to a research paper and is mapped to all the metrics associated with the research paper. Consequently, these outcomes are further considered while generating the Potential Evaluation Plan and the Potential UX Outcome.
In addition to the recommended metric outcomes, our system generates potential risks associated with the user system, using the AIAAIC repository. We extract the top three incidents that have a Euclidean distance of less than 0.5 by performing a vector-based similarity search, using FAISS, between the user system description, and the AI system description present in the AI Incident dataset. Subsequently, we extract risks associated with these AI systems, and provide them along with the user system description to a LLM, to filter the most relevant risks for the system. The risks identified are contextualized based on the initial system description, with corresponding URLs for related incidents also available to the user.
Once the metrics have been finalized by the user, along with the corresponding metric outcomes, a Potential Evaluation Plan and Potential UX Outcome is recommended to the user. The Potential Evaluation Plan includes detailed instructions, regarding the usage of each metric and the methods for measuring each metric, based on the user’s System Description and initial evaluation plan. The Potential Evaluation Plan is generated by providing the system description, a list of selected metrics, the specified metric outcome, and methods used to previously evaluate the UX Metrics, to a LLM, and subsequently, prompting the LLM to generate the Potential Evaluation Plan. The Potential UX Outcomes are generated by considering the list of selected metrics, desired results from the metric outcomes, initial system description, and also the risks associated with the system.
3.3. Technical Evaluation
We conducted a small-scale evaluation to compare EvAlignUX ’s understanding of research metrics with human understanding and assess the quality of its metric suggestions based on a given narration.
3.3.1. Evaluating LLM’s Ability to Extract the Metrics
To assess GPT’s capability in identifying evaluation metrics in research papers from a predefined list of candidate metrics, we adopted a similar strategy as (Pang et al., 2024) and randomly selected 15 papers from our initial pool of research articles. The evaluation metrics used in these papers were extracted using our GPT prompt. One author reviewed each identified metric and rated it from three aspects: whether the metric was actually measured, whether the metric was in the candidate list and whether the reasoning or citation provided makes sense. A second author verified these assessments. A metric was considered accurate if it fulfils requirement one and two. Common sources of inaccuracy included:
The metric identified was not actually measured in the evaluation part of the paper (i.e. hallucination (Ji et al., 2023)). There were three cases for this: a) GPT inferred metrics based on descriptions from the paper, but the metrics were not explicitly measured in the evaluation process. b) GPT synthesized metrics from the results, but the metrics were not explicitly measured in the evaluation process and c) GPT making recommendations for a metric that should have been used but was not used, based on the context of the paper. The case for (a) and (b) was often that GPT picked out a statement or a conclusion from a qualitative result described in the paper. However, such qualitative results were synthesized from interviews or surveys, and therefore the identified metric was not explicitly measured in the paper. For example, GPT identified ”attitude towards the agent” as an evaluation metric from (Wambsganss et al., 2020), quoting ”students made positive comments about the interaction with CAs related to the perceived level of enjoyment and interaction.” While attitude towards the agent was not explicitly measured in the paper, it was described in the findings and was thus misidentified. For (c), an example is GPT identifying ”privacy concern” as a metric from (Lee et al., 2021), quoting ”because chatbot use can result in users’ self-disclosure of sensitive topics” from the discussion section. Although it makes sense to measure ”privacy concern” based on the context of potential self-disclosure of sensitive information, the said metric was not explicitly measure in the paper so it was marked as inaccurate. The metric identified was outside the list of candidate metrics. There were two cases for this: a) GPT reworded the metric in the candidate list to express the same meaning, and b) GPT made up a metric that was not included in the candidate list. For (a), an example was GPT identifying ”utterance length” from (Thomas et al., 2020), while the actual metric in the candidate list is worded as ”length of utterance”. Regarding (b), ”transparency” was identified as a metric from (Pecune et al., 2019), while ”transparency” is not part of the candidate list.
We found that out of the 102 metrics identified in the 15 papers, 94% (96/102) were indeed measured, 92% (94/102) were in the candidate list and 88% were accurate (90/102), as there were two overlapping metrics across requirement one and two. 100% (102/102) of the reasoning and citations made sense. Overall, the results suggest that with our prompt, GPT’s ability to identify metrics is largely reliable.
3.3.2. Evaluation of Metric Recommendation
We evaluated our system using a sample of 40 papers as ground truth, including papers in AI, wearable technology, VR/AR, conversational agents, etc. Our system is supported by GPT-4o, and we also benchmarked our system against the GPT-4o LLM model to recommend UX metrics using precision, recall, and F1 score (Tamm et al., 2021; Fayyaz et al., 2020). The baseline GPT-4o and our system were given the same set of inputs, including the system description and the corresponding list of indexes. Then, each sample was processed three times by both our system and the baseline to generate a list of metrics. The results obtained from both the systems were generalized to reflect the naming standards used in test dataset. To perform the final comparision, we calculated the precision, recall, and f1 score for each subsequent run, and the final metric scores obtained were averaged to obtain the overall mean scores.
| Model | Mean Precision | Mean Recall | Mean F-1 Score |
| GPT-4o | 0.096 | 0.203 | 0.121 |
| EvAlignUX | 0.156 | 0.304 | 0.195 |
Table 1 demonstrates that our system significantly outperforms the baseline GPT-4o, with a 6% increase in precision, a 10.1% increase in recall, and a 7.4% increase in F1 score. We further conducted a pair-wise t-test by processing each sample 10 times on both the systems, to compare the metric mean between EvAlignUX and baseline (). The t-test demonstrate that their is there is a statistically significant difference between the performance (precision, recall, and F1 score) of the two models on the same dataset. Additionally, we observed that our system recommends more relevant UX metrics for a given system description than the baseline GPT-4o. For instance, when evaluating Tan et al.’s work (Tan et al., 2024), our system recommends using Perceived Naturalness to measure how naturally the AI-generated visuals integrate with the audiobook narrative. However, when evaluating the baseline GPT-4o, none of the generated metrics explicitly address the naturalness of the integration between visuals and audio. Furthermore, when testing the system on Wang et al (Wang et al., 2024), our system recommends evaluating Emotion Awareness to monitor the system’s ability to recognize and appropriately respond to users’ different emotional states, which is also measured in the research However, none of the metrics recommended by the baseline model address the evaluation of the varied emotional states of the user. Additionally, when using (Sun et al., 2024), the system recommends measuring Perceived Intelligence to assess how users perceive the system in generating coherent motions, which the baseline model does not explicitly recommend. The results indicate that our system outperforms the baseline GPT-4o in recommending UX metrics to users.
4. User Study
In this section, we outline how we evaluate EvAlignUX ’s user experience, and we aim to explore the following key research questions.
RQ1: How effectively does the use of EvAlignUX impact researchers in developing their UX evaluation plans?
RQ2: How do users leverage the EvAlignUX features to improve thought processes of developing UX evaluation plans?
RQ3: How does the users’ background impact their use of EvAlignUX ?
4.1. Participant
Our study was approved by the university’s IRB. We recruited 19 participants from social media. All participants reported they had extensive experience with HCI/UX research, 11 participants (57.9%) have 3-4 years experience, 7 participants (36.8%) have 1-2 years experience, only one participant has less than one year UX experience. Regarding participant’s AI usage in research study, the majority of participants used AI to help with simple tasks, such as coding & writing (17, 89.5%), literature review (7, 36.8%), and data analysis (11, 57.9%),. Notably, we found that few participants used AI for tasks that require rigor and creativity, including research ideaiton (7, 36.8%), design prototyping (7, 36.8%), research evaluation plan development (3, 15.8%), and usability testing (1, 5.3%). The background information of participants is shown in Appendix 6. Each user study lasted 1.5 to 2 hours. Participants received 20 USD per hour as compensation in the user study.
4.2. Procedure
Figure 6 demonstrates the study procedure. ① The study began with the background collection, including their UX experience levels. Researchers then introduced the study context and explained the tasks. Then, they used EvAlignUX ’s landing page to spend about 10 minutes entering their project description, evaluation plans, and expected outcomes. After that, they completed an initial survey that capture their perceptions of the proposed project plans (survey 1, initial plan). Then, participants used the Zoom Whiteboard to draw their thought processes for developing the UX evaluation plan and think-aloud to ensure their ideas were communicated clearly.
In EvAlignUX , participants were guided through the system features and asked to refine their project plans using all available features. After experimenting with the features and achieving satisfactory results, participants used the ”Export Data” button to review, regenerate, and re-edit their project plans. Upon finalizing their plans, they completed exit surveys (survey 2, revised plan; survey 3, EvAlignUX ) to evaluate their perceptions of the generated plans and their experience with the system. During the EvAlignUX process, the think-aloud method was employed to observe and record participants’ interactions and thoughts. Next, participants used Zoom Whiteboard to revise their thought process for developing the UX evaluation plan if they felt there were any changes. Finally, participants were interviewed to delve deeper into their overall experiences and gather additional insights. This step provided a comprehensive understanding of user perceptions and allowed for collecting feedback on both the proposed plans and the system features.
4.3. Measures
We recorded the following data to assess participants’ performance and perceptions during the study: task logs, thought process diagrams, think-aloud data, Likert-scale questionnaire responses, and interview transcripts.
4.3.1. System Logs
For system log data, we firstly collected user’s input to the system, including project description (i.e. What are the main objectives of your project?), UX evaluation plans (i.e. How do you plan to evaluate the UX of your project?), and expected outcomes (i.e. What outcomes do you expect from this evaluation?) as well as their updated ones. Additionally, we logged their usage behavior data, including (1) the number of times a user regenerate indexes, (2) the total number of metrics selected, (3) the number of metrics added through graph, (4) the number of outcomes clicked including both selected and deselected, and (5) the number of risks clicked. We also logged their revised system description, UX evaluation plans, and expected outcomes.
4.3.2. Thought Process Diagrams
We collected user’s thought process visualization for developing a UX evaluation plan using Zoom Whiteboard before and after using the EvAlignUX to see how AI-based influenced their thinking.
4.3.3. Questionnaire
The questionnaire is divided into three main parts:
Background information: Users answered questions related their past UX experiences, AI literacy, and AI usage for research purposes.
Perceptions of initial and revised UX study plan (survey 1 and 2): We collected user perceptions before and after they explored EvAlignUX to determine if there were significant differences between their perceptions of their UX research plans. The survey items are adapted from main schemes for analyzing UX evaluation methods (Vermeeren et al., 2010; Venable et al., 2012; Dean et al., 2006), including novelty, relevance, efficiency, completeness, clarity, concreteness, quality, feasibility, rigorous, risks considerations, and confidence.
System feature perceptions: For each metric, we asked participants to rate their perceptions of each feature across 8 Likert-scale items (survey 3): enjoyment of use (Davidson et al., 2023), intention to use it in the future (Everard and Galletta, 2005), perceived usefulness (Davis, 1989), trust in the suggestions and decisions provided (Everard and Galletta, 2005), perceived cognitive load in understanding the information (Paas et al., 2008), inspiration for exploring and developing UX research (Chu et al., 2010), the system’s contributions to the UX evaluation plan, and the system’s contributions to the UX outcome.
4.3.4. Think Aloud
We collected their think-aloud while using the system features as well as their thought process diagrams. For example, we asked to provide us with reasoning behind their decision making process.
4.3.5. Interview
Finally, we conducted a short follow-up semi-structured interview. We asked questions in three main categories with follow-up questions based on the interviewee’s responses.
Perceptions of features and their role in decision making for proposing a UX plan: Which feature is the most impressive? Which feature is the least impressive? How do these features support your decision making?
Overall user experience and potential improvements: After using the system, how useful do you feel about the system? Were there any features or information that you felt were lacking? How do you view your original and revised plans?
Follow-up questions about their research mental model: What differences do you notice between your previous research process and the process of interacting with EvAlignUX ? How do you perceive the strengths and weaknesses of these strategies?
4.4. Data Analysis
We analyze each participant’s ① thought process diagrams, ② quantitative data from surveys and logs, and ③ user decision making reasoning through think-aloud, interview transcripts, and UX research plans.
4.4.1. Thought Process Diagrams Analysis
We standardized thought process diagrams drawn by participants by coding each illustrated step. Modifications, such as envisioning AI assistance or adding new steps, were identified. These changes were further analyzed by examining their diagrams with the think-aloud data. We summarized the benefits participants identified where AI could assist in proposing a UX research plan.
4.4.2. Quantitative Data Analysis
We analyzed participants’ responses to the two post-task surveys. For all these measures, we conducted a Shapiro-Wilk test (González-Estrada and Cosmes, 2019) and found that all data were not normally distributed. Given the nature of our data, we applied Kendall’s tau correlation to explore relationships between user experience levels, system usage, and perceptions of changes in plans and system features(Bolboaca and Jäntschi, 2006; Kendall, 1938). We also reported effect sizes using standardized methodologies (Dixon and Massey Jr, 1951). All statistical analyses were conducted using Python 3.12.
4.4.3. Behavioral Analysis through Think-Aloud and Interviews
A thematic analysis (Braun and Clarke, 2012) was performed on data from think-aloud sessions and interviews. Two researchers collaboratively identified common themes through an iterative process. This analysis helped map how EvAlignUX features address identified challenges from the thought processes and outline the pros and cons of these features.
5. Findings
This section presents findings on how participants’ perceptions towards their proposed UX evaluation plans changed through the use of EvAlignUX (RQ1) in 5.1, their feature usages and perceptions (RQ2) in 5.2, and how these perceptions varied based on their UX research background (RQ2) in 5.3. Participant quotes are indicated by quotation marks, while system features are formatted in code style.
5.1. EvAlignUX Significantly Improving the UX Evaluation Plans (RQ1)
In this section, we compared participants’ ratings of their original and revised UX plans across 11 measurement criteria. Of these, 9 measurements showed significant improvement in the revised plans compared to the original. Five participants fully accepted the system-generated evaluation plan without modification. On average, 60.1% (SD=0.33) of the generated UX evaluation plan and 59.6% (SD=0.38) of the generated UX outcome were incorporated into their final evaluation plans. The results are shown in Figure 7, with detailed information available in Appendix Table 7.
Overall, the revised plans were more specific. The increase scores of Concreteness (W = 24.0, = .032∗, ), Completeness (W = 16.0, = .004∗∗, ), Rigor (W = 15.0, = .009∗∗, ), and Quality (W = 11.0, = 0.004∗∗, ) suggest that participants added more detailed information to their revised plan. Several participants only had a vague idea about which variables to measure (e.g., P7, P14, see Table 2). With EvAlignUX , they noted how it helped them improve the details in the UX plans. For example, P14 later strategically specified six measurements by questionnaires and system logs, with each in great detail; and P7 copied a holistic UX evaluation plan using a survey, system logs, and interview protocols, each with specific details added.
Participants not only added more details to their initial plans, but they also perceived their plan more concise and relevant since Clarity (W = 4.0, = 0.005∗∗, ), and Efficiency in achieving research goals (W = 12.0, = .012∗, ) also demonstrated a considerable improvement. Participants made meaningful revisions for clarity. For instance, P5 removed texts related to assessment from their initial system description, as they found it more suitable as a part of the evaluation plan instead of putting it into the system narration. Also, several participants praised the "Metrics list" feature for clarifying metric relevance, noting it showed ”how this metric can be relevant in the context (P3)” and ”encouraged the reasoning of using this metric (P17)”. P8, during a think-aloud session, explained, ”I just selected ‘information retention’ because my project is about memory, and the definition reminds me of the importance of evaluating how well my participants retain information to measure the effectiveness of counseling sessions in my project.”
| Input | ID | Initial Plan | Revised Plan |
| UX Plan | P7 | To conduct a semi-structured interview with experts to see how they collaborate with our system to analyze the video | Use a survey to measure the trust experts have in the AI’s analysis capabilities. Questions could focus on reliability, perceived competence … Discuss their perceptions of the assistant’s expertise, any concerns about the analysis, and instances where the assistant’s suggestions were particularly effective or ineffective. |
| UX Plan | P14 | Ask users to experience this tool and other similar ones and then report usability, satisfaction, and others with questionnaires. | Collection of short open-ended responses from participants to analyze qualitatively. Ask pre and post survey questions about their decision confidence (how accurate they feel about their prediction of what the AI will do), perception of how clear the AI visualization was, and how accurate their prediction was. Do these before and after each action. |
| UX Outcome | P13 | Low mental load, higher satisfaction, higher quality of subfield extraction. | I can expect the tool to provide accurate and useful summaries for both lay people and HCI researchers, leading to high user satisfaction and effective functionality. The interactive visualizations will likely be well-received for their ability to save time and allow quick exploration. |
| UX Outcome | P11 | It outperforms the baseline techniques in the education effect and subjective preferences. Achieved this results through a more engaging experience, which also lowered participants’ understanding difficulty. | (…Same with initial…) There are potential risks alongside the expectations that the research outcome may not align with the intended education effect and the system may not be generalizable. However, these may could be partially reflected in the objective metrics and be acknowledged in the limitations. |
Participants also felt more confident and believed their plans is feasible as Feasibility (W = 16.5, = 0.035∗, ) and Self-Confidence (W = 18.0, = .014∗, ) showed significantly increases. For example, participants felt they could know how to cite (P12, P15), ”measure (P1, P3, 14)”, ”design questions (P1, P9, P17)”, “learn previous findings (P8, P2)” for metrics using the metric feature, making the UX plan tangible, and even “how to collect the data specifically” (P13, see Table 2). They also gained confidence, especially when noticing prior work used the same metric selections. As P3 expressed, ”I’m really happy someone else also uses ’improvement in social skills’ in this context.” The validation further reinforced their self-confidence in the chosen metrics and plans.
Participants also significantly revised their plans to include more risk considerations (W = 10.0, = 0.007∗∗, ). For example, building upon original UX outcomes, P11 expanded, ”There alongside the expectations the research outcome may not align with education effect the system may not be generalizable. However, these may could be partially reflected in the objective metrics and be acknowledged in the limitations.” P11 took inspiration from concerns about generalizability from the system’s output (see Table 2, where ”Generalization Failure” was mentioned, stating, ”Thinking about the potential biases made me rethink how I approach user data to ensure that I don’t overlook important issues.”
However, there was no significant changes found in ”Novelty” (W = 22.0, = 0.083, ) and UX plans’ ”Relevancy” to research questions. Several participants felt the system was hard to change. P4 noted, ”The core idea is already set,” while P6 said, ”It’s challenging to introduce major changes.” Participants also reflected that ”It is hard to justify the novelty without referring to the actual paper and verify it on my own” (P4).
5.2. EvAlignUX Feature Use and Perceived Value (RQ2)
In this section, we discuss how the individual features of
EvAlignUX helped overcome different challenges in developing evaluation plans while also streamlining the thought process. In particular, three features, metrics graph, metrics list, and outcomes, are regarded as the most effective.
5.2.1. The Value of EvAlignUX Features in Enhancing UX Evaluation
We observed how the use of individual features and the combined use of different features of EvAlignUX enhances the user experience. In Figure 8, we summarized a group of prototypical questions that UX researchers asked themselves when developing their UX evaluation plans, annotated from think-aloud data during participants’ thought processes, which highlighted five perceived benefits of AI: justify, inform, support self-critique, streamline, and vision.
Participants mentioned how AI could be used to Inform them knowledge in prior literature, e.g., to quickly “suggest relevant literature” (P15, 18), “summarize existing evaluation methods based on the research context” (P3, P11), supporting researchers in understanding appropriate metrics by providing “definitions” (P8), “citations” (P3), “how it can be used” (P16), and “constructs adopted” (P17). For example, P19 combined the Index and Metrics List features. P9 initially felt that EvAlignUX did not accurately capture their intention. They proposed a version control system, but EvAlignUX recommended metrics that were more likely to measure ”machine learning efficiency,” such as accuracy. However, after changing a few indexes (e.g., adding two design novelties: graph-based version control, and automated design documentation), they found EvAlignUX suggested new metrics, such as interaction experience, overall satisfaction, perceived usability, and task success, to be much more accurate and aligned with their research.
Many participants recognized the value of using AI to visualize literature networks, which help them better Justify metric choices to align with research goals. This capability was particularly appreciated as it facilitates “metric selection across various venues and disciplines” (P17), “highlight connections between metrics” (P16) to determine whether two metrics are “interrelated or distinct” (P13) and support decisions depends on “popularity and relevance in the field” (P9) in a way that “aligns with their (my) research goal” (P6). For example, P9 used Metrics Graph and Outcome together. P9 designed a conversational agent for judges. After reviewing the graph, they found that user satisfaction, accuracy, and task completion were connected, which motivated him to think about the relationship between these three metrics. However, he found that while certain metrics aligned with their research goals, not all findings were applicable to their specialized context. They noted, no emotions or diversity should be included as my evaluation outcomes; otherwise, the judges will be easily biased.
Participants recognized the value of using AI to help them better Self-Critique during their UX evaluation plan development process. This capability was particularly appreciated as it facilitates “prompt feedback” (P1) with “new perspectives” (P15), acting as a “teammate” (P6, P16) or “partner” (P18) in scholarly settings, to comment on their proposed plans are “effectiveness and limitations” (P13). One participant also mentioned about its potential to support ”overcoming social inhibitions, like hesitancy in seeking advice or posing questions to peers too frequently” (P16). As P4 commented that, ”It’s nice also to have EvAlignUX where you can throw ideas around, and then it will guide you directly in your research project, what to think about—not just a chat with somebody, which can sometimes be embarrassing.”
Participants also recognized the value of using AI to reduce cognitive load, which helps them better Streamline their UX evaluation development to align with dynamic project timelines. This capability was particularly appreciated as it ”simplify decision-making processes in UX design” (P16) to determine whether new updates are ”directly contributing to the UX” (P13) of their original research goal or not. For example, P4 said, ”Generally if I need to propose a UX plan, I probably need to spend like two days at least to collect all the paper. I believe your system can actually help us [HCI researchers] to shorten the time. Because every paper here is coming from valid sources, so we can trust the result.” Furthermore, participants also shared the value of using the system for re-structuring their thought-process of the UX plan development. Figure 9 is an example illustrating how P2 modified the literature review process by using AI to conduct the literature review and brainstorm metrics earlier in the process.
Nearly all participants recognized the value of using AI to promote early detection of risks, which helps them better Vision their UX planning. This capability was particularly appreciated as it facilitates “spotting potential risks before they turn into real headaches after the study is done” (P5), and supports decisions based on “minimizing unexpected challenges”(P9). As P10 said mentioning the Risk feature, ”It doesn’t simply show a list of risks but encourages users to click the associated link to review the actual report.”
5.2.2. Enriched Thought-Processes with the Support of EvAlignUX
Besides how participants used the EvAlignUX features differently, they also presented diverse thought processes during their development of UX evaluation plans. When they updated their thought processes, they all revised their literature review process by incorporating relevant EvAlignUX features. More than half of the participants expressed that AI facilitates their iterative loops during developing an evaluation plan (8 participants). For instance, Figure 10. illustrates P13, a UX researcher with 1-2 years experience, revised thought process after using EvAlignUX , showing AI’s role as a supportive teammate in this process. The AI-suggested metrics, as highlighted by the purple boxes, prompted participants to critically assess the relevance and appropriateness of AI’s recommendations, ensuring alignment with system design goals and expected results. This iterative reflection, where P13 evaluate the added value of AI’s input, emphasizes AI’s role in fostering critical thinking and refinement.
Similarly, Figure 11. demonstrates how P11, with 3-4 years of UX experience, views AI can benefit the process of the iterative metric selection process.
We also observed that UX risks and outcomes were added into the thought process, regardless of participants’ experience level. Overall, 42% of participants incorporated new components into their UX evaluation development thought processes after using EvAlignUX . Notably, these included ”Propose measurements” (3 junior UX-level participants, mixed research stages), ”UX risks” (mixed UX levels, mixed research stages), and ”Research outcomes” (mixed UX levels, mixed research stages). These additions reflect a newly gained awareness of potential challenges and impacts in their UX evaluation plan development. For example, Figure 12 shows how P7, a UX researcher with 1-2 years of experience, added two new considerations: defining the ideal outcome of the evaluation and identifying potential risks associated with the chosen measurements to their thought process after using EvAlignUX .
5.2.3. The Significant Use of Hypothesized Outcome and Metrics Exploration Features
Based on these observations, we further analyzed the relationship between users’ system log variables (e.g., number of times indexes were regenerated, number of metrics selected via the Metrics List, number of metrics selected or deselected via Metric Graphs, number of outcomes clicked, number of risks clicked) and their perceptions of five different features across 11 perception variables (e.g., enjoyment, intention, usefulness, cognitive load, inspiration), see Section 4.3.3 for measurements. A total of 55 pairs were examined, with 6 showing significant results. All significant results are presented in Table 3.
| Dependent Variable | |||||
| Feature Usage Log | Feature: Perception Scores | Mean | SD | 95% CI | |
| # of metrics added by graph | Metrics Graph: Cognitive Load ∗ | 2.74 | 1.15 | -0.40 | [-0.72, -0.07] |
| # of selected results | Outcome: Enjoyment ∗∗ | 3.32 | 1.16 | 0.57 | [0.28, 0.82] |
| Outcome: Intention to Use ∗ | 3.63 | 1.07 | 0.40 | [0.05, 0.66] | |
| Outcome: Contribution to UX evaluation∗ | 3.42 | 1.39 | 0.44 | [0.08, 0.71] | |
| Outcome: Contribution to UX outcomes∗∗ | 3.37 | 1.50 | 0.49 | [0.16, 0.75] | |
The result showed that selecting more potential outcome items (i.e., previous research findings) is associated with higher user ratings of the Outcome feature, both enjoyment (r = 0.57, 95% CI [0.28, 0.82], p ¡ 0.01∗∗) and in future use (i.e., intention to use: r = 0.40, 95% CI [0.05, 0.66], p ¡ 0.05∗). Additionally, users rated this feature more positively in terms of its contribution to revising their project idea (r = 0.45, 95% CI [0.05, 0.76], p ¡ 0.05∗), UX evaluation plan (r = 0.44, 95% CI [0.08, 0.71], p ¡ 0.05∗), and UX outcomes (r = 0.49, 95% CI [0.16, 0.75], p ¡ 0.05∗). The findings from relevant HCI domains encouraged them to think about questions such as what demographic is left in the original plan? (P5) What factor was not considered yet? (P14) These thoughts made them to refer back to their original project idea, UX evaluation plan, and UX outcomes. P4 selected 15 potential outcomes that were derived from prior work, they commented that for instance, ”measuring trust and confidence in the chatbot’s abilities and context understanding” was ”exactly what they were looking for and this point really hit the point”. P4 further commented, they liked this feature because they felt ”the generation understands my initial thoughts about the project… It helps me finalize the details of my plan.”
Additionally, the results also indicated that adding more metrics to the graph was associated with a reduction in perceived cognitive load (r = 0.40, 95% CI [0.07, 0.72], p ¡ 0.05∗). Participants found using the feature of Metrics Graph simple – not only showing ”an aerial view of how these metrics tie together” (P6) but also the usage of each metric in prior research, which is ”easy to understand” (P8).
5.3. EvAlignUX Perception Varying by Research Background(RQ3)
In this section, we examined the relationship between users’ backgrounds and their perceptions of 5 different features across 9 perception variables (e.g., enjoyment, intention, usefulness, cognitive load, inspiration, see measurements in section 4.3.3), aiming to identify the users who benefit the most from our system and the users we still need to improve support, with 7 showing significant results. All significant results are shown in Table 4.
| Dependent Variable | |||||
| Background | Feature: Perception Scores | Mean | SD | 95% CI | |
| UX experience | Index: Perceived inspiration ∗ | 4.00 | 0.94 | -0.44 | [-0.73, -0.05] |
| Metrics list: Perceived inspiration ∗∗ | 4.11 | 0.94 | -0.67 | [-0.84, -0.43] | |
| Metrics Graph: Perceived inspiration ∗ | 3.89 | 1.10 | -0.49 | [-0.77, -0.13] | |
| Outcome: Perceived inspiration (n.s.) | 3.68 | 1.00 | |||
| Risk: Perceived inspiration∗ | 3.68 | 1.33 | -0.42 | [-0.72, -0.04] | |
5.3.1. The Extent of Perceived Inspiration Varies by Users’ Research Experience
Overall, the perceived value of providing inspiration through EvAlignUX was significantly negatively correlated with participants’ background, with the exception of the Outcome feature. Namely, participants with less UX experience provided higher scores for perceived inspiration, except with the Outcome feature. Despite this, all participants rated the inspirational measure highly (M = 3.58–4.00, SD = 0.94–1.33).
For Index, some participants with less experience shared that, the feature helps to structure research ideas using relevant terminologies as sometimes they ”just have a vague thought without knowing the exact terms” (P9). P6, with 1 year of UX experience, shared that the feature is a ”Very good at pinpointing important aspects, like user-centered design and stakeholders (researchers). It did a good job of checking what the user intended. I wouldn’t say anything was wrong or needs changing. Everything was spot on. There are certain things you can add, but they’re already in line with what was checked before.” For experienced researchers like P2, she mentioned that the feature helps to distinguish ”umbrella terms” when they only have a vague understanding of what they are measured: ”It lowered the effort to identify the literature that is actually related. For instance, I saw three terms, knowledge acquisition, knowledge retention, and knowledgability. If I just use the wrong term, I may think that no one has done it before, but it turns out people have already done it.” While experienced users acknowledged its utility for combining suggested indexes in their paper searches, they felt it lacked novelty, as many of the indexes were already familiar (P2), ”I am looking for something that I have no knowledge about.” Instead, they tend to ”self-validation whether they are on the right track” (P4) instead of seeking inspiration for relevant knowledge.
For Metrics List and Metrics Graph, less experienced users valued it for revealing ”metrics [they] hadn’t considered” (P7). As P3 (1-2 years of UX experience) mentioned, ”Because generally if I need to do that (on my own), I probably need to spend like 2 days at least to collect all the paper, which is really bad. I think EvAlignUX can actually help us to shorten the time. Because every paper here is coming from valid sources, such as top conferences, so we can trust the result.” Similarly, P6 said, the graph gives her serendipity, ”An aerial view of how these evaluation schemes tie together, like from usefulness to ease of use, helps when writing the paper or an evaluation scheme. Certain things in usefulness and trust are often correlated, as shown by the thickness of the line, which is inspiring.” The graphs help researchers to identify is their chosen metrics are “novel”, because the size of the nodes represent the frequency a metric is used in a research domain; this also help them to discover some “under-studied” metric in a domain; also for commonly used metric, they got to know what are the common practices. As P7 said, ”I feel like changing my selections based on the graph resulted in better and more related contributions.”
This pattern extended to the Risk feature, where P6 with 1-2 UX research experience found it interesting to identify ”risks that directly related to the system.” More experienced researchers, like P1, mentioned that ”The risks mentioned here are not hardlines, meaning that if users are not followed, then the system will fail. For instance, if the evaluation plan doesn’t get the IRB approval, they cannot continue to the next step; whereas the potential risk information here is more likely to be soft suggestions, which may not contribute to actual revision.”
Interestingly, inspiration from the Outcome feature did not vary with UX experience, aligning with participants drawing of thought-process, where outcomes were often missing from initial UX plans regardless of participants’ UX experience. The overall high mean score of rated inspiration from Outcome suggests the feature resonated across all users, though for different reasons based on experience level. Several participants with less UX experience found the Outcome feature useful for refining their research plans, noting how it helped them ”rephrase [their] outcome to make it more detailed” (P13). P7 said, “I haven’t thought about it before, but now that I saw it here, that can be an useful contribution…It’s really cool that they come from actual papers and aren’t fabricated or made up.” Whereas, P4, with 3-4 years of UX experience, noted that “without going into too much like spending a day thinking about it (UX outcomes). I think that it (your suggested UX outcomes) is a really great restatement of my original one.”
6. Discussion
In this section, we explore methods for promoting a mindset-focused UX evaluation in HCI education and practice, examine the potential risks of using large language models for evaluation, and discuss potential design implications yielded from our findings.
6.1. From Method-Centric to Mindset-Centric: a Paradigm Shift of Creating UX Evaluation Plans
In this study, we found that highlighting UX metrics as a center of UX evaluation empowers UX researchers holistically, enhancing both their perceived quality of their proposed plans and their confidence in their UX evaluation plans (RQ1). Furthermore, the combined use of different EvAlignUX components for refining their UX plans allowed participants to self-validate, critique, and rationalize their evaluation choices (RQ2), with their perceived inspirations from the system varying according to their UX experience (RQ3). Meanwhile, we observed their thought processes evolving from asking general questions to using AI to support iterative reasoning for their chosen metrics, considering various types of information, including previously not considered aspects, such as outcomes and risks.
UX evaluation plays a vital role in ensuring the quality of HCI system research, and it is a growing field with diverse approaches (Pettersson et al., 2018). Although several UX evaluation methods have been outlined in previous literature reviews (Pettersson et al., 2018; Vermeeren et al., 2010; Rivero and Conte, 2017; Darin et al., 2019; Kieffer et al., 2019), few interaction designs have been proposed to support UX researchers in developing comprehensive UX evaluation plans. Gray’s previous work (Gray, 2016) found that many UX designers and practitioners view UX as a set of rigid methods rather than a ”mindset,” suggesting that HCI education should move beyond teaching predefined methods and instead focus on fostering the adaptability and improvisation skills necessary for UX designers to address the unique challenges they face. A mindset approach to UX evaluation, as emphasized in our findings, aligns with their argument.
Several scholars emphasize adaptability and context-sensitive practices in UX research. Vermeeren et al. (Vermeeren et al., 2010) note that no single UX evaluation method suits all projects, urging researchers to adapt methods to specific contexts. Rivero et al. (Rivero and Conte, 2017) advocate for iterative frameworks that incorporate feedback, while Darin et al. (Darin et al., 2019) stress the importance of critical reflection throughout the evaluation process. Building on these insights, our study highlights the role of metrics in establishing a mindset-oriented approach to UX evaluation. This shift is essential for advancing the field and ensuring UX evaluations meaningfully contribute to the success of HCI systems research.
6.2. Design Implications: Using LLM for UX for Interdisciplinary and Crossed-Studies Research
Our findings reveal that EvAlignUX provides five key benefits for developing UX research plans. However, we observed varying usage patterns among participants (RQ1, RQ2): some enriched system-generated plans with detailed context, others focused on applying specific metrics, and some simply copied the plan without critical evaluation. We found that these patterns were, to some extent, linked to participants’ UX experience (RQ3).
While some participants viewed EvAlignUX as a potential “teammate” in developing UX plans, the challenge remains in ensuring that LLM usage does not lead to the deskilling of UX practitioners’ research expertise. Previous literature has highlighted the ethical concerns and potential harms of over-reliance on LLMs in various domains. For instance, Ahmad (Ahmad et al., 2023b) found that excessive dependence on AI in education led to students becoming passive, diminishing their decision-making abilities and negatively impacting cognitive skills. Similarly, in human-AI co-creative writing, AI outputs are heavily influenced by the distribution of their training datasets (Roselli et al., 2019; Argyle et al., 2023). Without considering users’ unique needs, over-reliance on these outputs can result in research homogeneity and a lack of creativity (Lee et al., 2024a; Rezwana and Maher, 2023). Additionally, issues such as hallucination, lack of transparency, and algorithmic bias have been identified as major barriers to effective AI collaboration (Zhai et al., 2024). Given the limitations of LLMs, future studies need to carefully consider and address the ethical concerns raised to support the critical thinking of UX researchers. Here are some potential research directions:
6.2.1. Mitigating Potential Over-Reliance and Triggering Deeper Thought Processes
One potential direction is to use AI systems that provide hints, or to use Question Bank (see Fig. 8), this approach may aim to augment researchers’ reasoning by prompting reflection and encouraging deeper engagement with their work. Essential features include controllability, and explainability, alongside clear disclaimers that warn users about the risks of over-reliance on generative AI (Liao et al., 2020b). Previous research highlights that transparent systems with clear AI outputs foster better human-AI collaboration by encouraging users to engage more critically with the tool (Wang et al., 2020; Glinka and Müller-Birn, 2023). Furthermore, studies indicate that embedding mechanisms that encourage active thinking—such as presenting alternative viewpoints or highlighting uncertainties—can help mitigate the risk of over-dependence on AI-generated solutions and foster deeper cognitive engagement (Bansal et al., 2021).
6.2.2. Using AI to Support Self-Critique
Our findings suggest that AI may assist in addressing challenges through self-critique mechanisms. For example, EvAlignUX uses its literature base to provide users with relevant information that encourages self-critique. However, unlike human experts, AI systems often fail to identify nuanced flaws in research plans that arise from researchers’ misunderstandings. Previous studies have shown that AI’s capacity for self-critique is limited, particularly when dealing with complex, contextual errors that require domain-specific knowledge or critical thinking (Bansal et al., 2021; Wang et al., 2021). AI’s ability to critique existing research, while promising, lacks the contextual depth found in human-to-human collaboration (Wang et al., 2021; Xu and Mulligan, 2022). Further research is needed to explore how AI can more closely mimic human critical evaluation in the research process.
6.2.3. Using AI to Mitigate UX Risks
A key design implication often overlooked is the concept of “impact visioning.” While only one participant in our study considered the potential risks during the initial development of their UX plans, previous studies have shown that negative consequences of new technologies can often be predicted in early design stages (Lee et al., 2024b; Pang et al., 2024). AI systems, by drawing on cross-disciplinary insights, may also help researchers anticipate and mitigate potential risks in UX evaluation, raising awareness of emerging concerns in technology use and design. Prior studies in AI ethics and risk management suggest that AI can be instrumental in identifying unintended consequences and risks in various domains, including UX design, by leveraging data across disciplines to foresee issues that might not be immediately apparent (Vinuesa et al., 2020; Floridi and Chiriatti, 2021).
6.3. Limitations and Future Work
We acknowledge several limitations that may affect both EvAlignUX ’s performance and the findings of our research. First, as an exploratory system, EvAlignUX incorporates a limited selection of papers, primarily focused on human-AI interaction. Participants proposing studies in interdisciplinary fields may find the system’s suggestions less relevant or accurate. Second, although we observed differences in participants’ perceptions and usage behaviors, the majority of participants were doctoral students, which may limit the generalizability of our findings. Third, for the EvAlignUX backend, the recommendation is confined to metrics that have been predefined in our database. Future studies should explore the continued expansion of the database to include a broader spectrum of metrics and online publication sources.
7. Conclusion
User experience is vital to any Human-Computer Interaction research, yet many UX designers and practitioners often rely on rigid evaluation methods rather than adopting a holistic approach. In this paper, we introduced EvAlignUX , an LLM-based tool designed to support UX evaluation development. Through an evaluation with 19 HCI researchers, we found that centering UX metrics in the evaluation process fostered a mindset that balanced holistic UX impact with contextual factors, evaluation practices, and research value. This approach significantly improved participants’ perceived plan quality and confidence, while also prompting their thought processes to consider risks when evaluating systems. For practical implications, we offer a Question Bank to guide future UX evaluation efforts; and suggest that future researchers can enhance EvAlignUX by incorporating domain-specific literature to further refine their evaluation planning. In a world where experience defines impact, a UX-driven mindset serves as the key to meaningful and lasting designs.
Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. 2119589. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Additionally, results presented in this paper were obtained using CloudBank (Norman et al., 2021), which is supported by the National Science Foundation under award No. 1925001.References
- (1)
- Abbasian et al. (2024) Mahyar Abbasian, Elahe Khatibi, Iman Azimi, David Oniani, Zahra Shakeri Hossein Abad, Alexander Thieme, Ram Sriram, Zhongqi Yang, Yanshan Wang, Bryant Lin, et al. 2024. Foundation metrics for evaluating effectiveness of healthcare conversations powered by generative AI. NPJ Digital Medicine 7, 1 (2024), 82.
- Ahmad et al. (2023a) Khlood Ahmad, Mohamed Abdelrazek, Chetan Arora, Muneera Bano, and John Grundy. 2023a. Requirements practices and gaps when engineering human-centered Artificial Intelligence systems. Applied Soft Computing 143 (2023), 110421.
- Ahmad et al. (2023b) Sayed Fayaz Ahmad, Heesup Han, Muhammad Mansoor Alam, Mohd Rehmat, Muhammad Irshad, Marcelo Arraño-Muñoz, Antonio Ariza-Montes, et al. 2023b. Impact of artificial intelligence on human loss in decision making, laziness and safety in education. Humanities and Social Sciences Communications 10, 1 (2023), 1–14.
- Ali et al. (2020) Zafar Ali, Guilin Qi, Pavlos Kefalas, Waheed Ahmad Abro, and Bahadar Ali. 2020. A graph-based taxonomy of citation recommendation models. Artificial Intelligence Review 53 (2020), 5217–5260.
- Argyle et al. (2023) Lisa P. Argyle, Ethan C. Busby, Nancy Fulda, Joshua R. Gubler, Christopher Rytting, and David Wingate. 2023. Out of One, Many: Using Language Models to Simulate Human Samples. Political Analysis 31, 3 (2023), 337–351. https://doi.org/10.1017/pan.2023.2
- Bangor et al. (2008) Aaron Bangor, Philip T Kortum, and James T Miller. 2008. An empirical evaluation of the system usability scale. Intl. Journal of Human–Computer Interaction 24, 6 (2008), 574–594.
- Bansal et al. (2021) Gagan Bansal, Besmira Nushi, Ece Kamar, Walter S Lasecki, Daniel S Weld, and Eric Horvitz. 2021. Does the whole exceed its parts? The effect of AI explanations on complementary team performance. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–16.
- Blondel et al. (2008) Vincent D Blondel, Jean-Loup Guillaume, Renaud Lambiotte, and Etienne Lefebvre. 2008. Fast unfolding of communities in large networks. Journal of statistical mechanics: theory and experiment 2008, 10 (2008), P10008.
- Bolboaca and Jäntschi (2006) Sorana-Daniela Bolboaca and Lorentz Jäntschi. 2006. Pearson versus Spearman, Kendall’s tau correlation analysis on structure-activity relationships of biologic active compounds. Leonardo Journal of Sciences 5, 9 (2006), 179–200.
- Bossen et al. (2016) Claus Bossen, Christian Dindler, and Ole Sejer Iversen. 2016. Evaluation in participatory design: a literature survey. In Proceedings of the 14th Participatory Design Conference: Full papers-Volume 1. 151–160.
- Braun and Clarke (2012) Virginia Braun and Victoria Clarke. 2012. Thematic analysis. American Psychological Association.
- Chaudhry (2024) Beenish Moalla Chaudhry. 2024. Concerns and Challenges of AI Tools in the UI/UX Design Process: A Cross-Sectional Survey. In Extended Abstracts of the 2024 CHI Conference on Human Factors in Computing Systems (CHI EA ’24). Association for Computing Machinery, New York, NY, USA, Article 83, 6 pages. https://doi.org/10.1145/3613905.3650878
- Chu et al. (2010) Hui-Chun Chu, Gwo-Jen Hwang, Chin-Chung Tsai, and Judy CR Tseng. 2010. A two-tier test approach to developing location-aware mobile learning systems for natural science courses. Computers & Education 55, 4 (2010), 1618–1627.
- Darin et al. (2019) Ticianne Darin, Bianca Coelho, and Bosco Borges. 2019. Which Instrument should I use? Supporting decision-making about the evaluation of user experience. In Design, User Experience, and Usability. Practice and Case Studies: 8th International Conference, DUXU 2019, Held as Part of the 21st HCI International Conference, HCII 2019, Orlando, FL, USA, July 26–31, 2019, Proceedings, Part IV 21. Springer, 49–67.
- Davidson et al. (2023) Shayn S Davidson, Joseph R Keebler, Tianxin Zhang, Barbara Chaparro, James Szalma, and Christina M Frederick. 2023. The development and validation of a universal enjoyment measure: The enjoy scale. Current Psychology 42, 21 (2023), 17733–17745.
- Davis (1989) Fred D Davis. 1989. Perceived usefulness, perceived ease of use, and user acceptance of information technology. MIS quarterly (1989), 319–340.
- De Meo et al. (2011) Pasquale De Meo, Emilio Ferrara, Giacomo Fiumara, and Alessandro Provetti. 2011. Generalized louvain method for community detection in large networks. In 2011 11th international conference on intelligent systems design and applications. IEEE, 88–93.
- Dean et al. (2006) Douglas L Dean, Jill Hender, Tom Rodgers, and Eric Santanen. 2006. Identifying good ideas: constructs and scales for idea evaluation. Journal of Association for Information Systems 7, 10 (2006), 646–699.
- Desmond et al. (2024) Michael Desmond, Zahra Ashktorab, Qian Pan, Casey Dugan, and James M Johnson. 2024. EvaluLLM: LLM assisted evaluation of generative outputs. In Companion Proceedings of the 29th International Conference on Intelligent User Interfaces. 30–32.
- Dixon and Massey Jr (1951) Wilfrid J Dixon and Frank J Massey Jr. 1951. Introduction to statistical analysis. (1951).
- Doshi and Hauser (2024) Anil R Doshi and Oliver P Hauser. 2024. Generative AI enhances individual creativity but reduces the collective diversity of novel content. Science Advances 10, 28 (2024), eadn5290.
- Edge et al. (2024) Darren Edge, Ha Trinh, Newman Cheng, Joshua Bradley, Alex Chao, Apurva Mody, Steven Truitt, and Jonathan Larson. 2024. From local to global: A graph rag approach to query-focused summarization. arXiv preprint arXiv:2404.16130 (2024).
- Everard and Galletta (2005) Andrea Everard and Dennis F Galletta. 2005. How presentation flaws affect perceived site quality, trust, and intention to purchase from an online store. Journal of management information systems 22, 3 (2005), 56–95.
- Fan et al. (2022) Mingming Fan, Xianyou Yang, TszTung Yu, Q Vera Liao, and Jian Zhao. 2022. Human-ai collaboration for UX evaluation: effects of explanation and synchronization. Proceedings of the ACM on Human-Computer Interaction 6, CSCW1 (2022), 1–32.
- Faruk et al. (2024a) Lawal Ibrahim Dutsinma Faruk, Mohammad Dawood Babakerkhell, Pornchai Mongkolnam, Vithida Chongsuphajaisiddhi, Suree Funilkul, and Debajyoti Pal. 2024a. A review of subjective scales measuring the user experience of voice assistants. IEEE Access (2024).
- Faruk et al. (2024b) Lawal Ibrahim Dutsinma Faruk, Debajyoti Pal, Suree Funilkul, Thinagaran Perumal, and Pornchai Mongkolnam. 2024b. Introducing CASUX: A Standardized Scale for Measuring the User Experience of Artificial Intelligence Based Conversational Agents. International Journal of Human–Computer Interaction (2024), 1–25.
- Fayyaz et al. (2020) Zeshan Fayyaz, Mahsa Ebrahimian, Dina Nawara, Ahmed Ibrahim, and Rasha Kashef. 2020. Recommendation systems: Algorithms, challenges, metrics, and business opportunities. applied sciences 10, 21 (2020), 7748.
- Floridi and Chiriatti (2021) Luciano Floridi and Mauro Chiriatti. 2021. Ethical challenges of AI in biomedicine. AI in medicine 84 (2021), 1–5.
- Fortunato (2010) Santo Fortunato. 2010. Community detection in graphs. Physics reports 486, 3-5 (2010), 75–174.
- Glickman and Zhang (2024) Mark Glickman and Yi Zhang. 2024. AI and generative AI for research discovery and summarization. Harvard Data Science Review 6, 2 (2024).
- Glinka and Müller-Birn (2023) Katrin Glinka and Claudia Müller-Birn. 2023. Critical-Reflective Human-AI Collaboration: Exploring Computational Tools for Art Historical Image Retrieval. Proceedings of the ACM on Human-Computer Interaction 7, CSCW2 (2023), 1–33.
- González-Estrada and Cosmes (2019) Elizabeth González-Estrada and Waldenia Cosmes. 2019. Shapiro–Wilk test for skew normal distributions based on data transformations. Journal of Statistical Computation and Simulation 89, 17 (2019), 3258–3272.
- Gray (2016) Colin M Gray. 2016. ” It’s More of a Mindset Than a Method” UX Practitioners’ Conception of Design Methods. In Proceedings of the 2016 CHI conference on human factors in computing systems. 4044–4055.
- Hart (2006) Sandra G Hart. 2006. NASA-task load index (NASA-TLX); 20 years later. In Proceedings of the human factors and ergonomics society annual meeting, Vol. 50. Sage publications Sage CA: Los Angeles, CA, 904–908.
- Hillman et al. (2023) Serena Hillman, Samira Jain, Craig M Macdonald, Elizabeth F Churchill, Carolyn Pang, Jofish Kaye, and Erick Oduor. 2023. Understanding and Evaluating UX Outcomes at Scale. In Companion Publication of the 2023 Conference on Computer Supported Cooperative Work and Social Computing. 466–469.
- Hyland et al. (1989) Michael E Hyland, Sidney H Irvine, Clive Thacker, Peter L Dann, and Ian Dennis. 1989. Psychometric analysis of the Stunkard-Messick Eating Questionnaire (SMEQ) and comparison with the Dutch Eating Behavior Questionnaire (DEBQ). Current Psychology 8, 3 (1989), 228–233.
- Inan Nur et al. (2021) Aulia Inan Nur, Harry B. Santoso, and Panca O. Hadi Putra. 2021. The method and metric of user experience evaluation: a systematic literature review. In Proceedings of the 2021 10th International Conference on Software and Computer Applications. 307–317.
- Ji et al. (2023) Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Ye Jin Bang, Andrea Madotto, and Pascale Fung. 2023. Survey of hallucination in natural language generation. Comput. Surveys 55, 12 (2023), 1–38.
- Kejriwal (2022) Mayank Kejriwal. 2022. Knowledge graphs: A practical review of the research landscape. Information 13, 4 (2022), 161.
- Kendall (1938) Maurice G Kendall. 1938. A new measure of rank correlation. Biometrika 30, 1-2 (1938), 81–93.
- Kieffer et al. (2019) Suzanne Kieffer, Luka Rukonic, Vincent Kervyn de Meerendré, and Jean Vanderdonckt. 2019. Specification of a UX Process Reference Model towards the Strategic Planning of UX Activities.. In VISIGRAPP (2: HUCAPP). 74–85.
- Kim et al. (2024) Tae Soo Kim, Yoonjoo Lee, Jamin Shin, Young-Ho Kim, and Juho Kim. 2024. Evallm: Interactive evaluation of large language model prompts on user-defined criteria. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–21.
- Kuang et al. (2024) Emily Kuang, Minghao Li, Mingming Fan, and Kristen Shinohara. 2024. Enhancing UX Evaluation Through Collaboration with Conversational AI Assistants: Effects of Proactive Dialogue and Timing. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–16.
- Lazar et al. (2017) Jonathan Lazar, Jinjuan Heidi Feng, and Harry Hochheiser. 2017. Research methods in human-computer interaction. Morgan Kaufmann.
- Ledo et al. (2018) David Ledo, Steven Houben, Jo Vermeulen, Nicolai Marquardt, Lora Oehlberg, and Saul Greenberg. 2018. Evaluation strategies for HCI toolkit research. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. 1–17.
- Lee et al. (2024b) Hao-Ping Lee, Yu-Ju Yang, Thomas Serban Von Davier, Jodi Forlizzi, and Sauvik Das. 2024b. Deepfakes, Phrenology, Surveillance, and More! A Taxonomy of AI Privacy Risks. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–19.
- Lee et al. (2024a) Messi H.J. Lee, Jacob M. Montgomery, and Calvin K. Lai. 2024a. Large Language Models Portray Socially Subordinate Groups as More Homogeneous, Consistent with a Bias Observed in Humans. In Proceedings of the 2024 ACM Conference on Fairness, Accountability, and Transparency (Rio de Janeiro, Brazil) (FAccT ’24). Association for Computing Machinery, New York, NY, USA, 1321–1340. https://doi.org/10.1145/3630106.3658975
- Lee et al. (2021) Yi-Chieh Lee, Naomi Yamashita, and Yun Huang. 2021. Exploring the effects of incorporating human experts to deliver journaling guidance through a chatbot. Proceedings of the ACM on Human-Computer Interaction 5, CSCW1 (2021), 1–27.
- Lewis (1992) James R Lewis. 1992. Psychometric evaluation of the post-study system usability questionnaire: The PSSUQ. In Proceedings of the human factors society annual meeting, Vol. 36. Sage Publications Sage CA: Los Angeles, CA, 1259–1260.
- Liao et al. (2020a) Q. Vera Liao, Daniel Gruen, and Sarah Miller. 2020a. Questioning the AI: Informing Design Practices for Explainable AI User Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3313831.3376590
- Liao et al. (2020b) Q Vera Liao, Daniel Gruen, and Sarah Miller. 2020b. Questioning the AI: informing design practices for explainable AI user experiences. In Proceedings of the 2020 CHI conference on human factors in computing systems. 1–15.
- Liao et al. (2024) Q Vera Liao, Mihaela Vorvoreanu, Hari Subramonyam, and Lauren Wilcox. 2024. UX Matters: The Critical Role of UX in Responsible AI. Interactions 31, 4 (2024), 22–27.
- Liu et al. (2024) Yiren Liu, Si Chen, Haocong Cheng, Mengxia Yu, Xiao Ran, Andrew Mo, Yiliu Tang, and Yun Huang. 2024. How ai processing delays foster creativity: Exploring research question co-creation with an llm-based agent. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–25.
- Marquardt et al. (2017) Nicolai Marquardt, Steven Houben, Michel Beaudouin-Lafon, and Andrew D Wilson. 2017. HCITools: Strategies and best practices for designing, evaluating and sharing technical HCI toolkits. In Proceedings of the 2017 CHI Conference Extended Abstracts on Human Factors in Computing Systems. 624–627.
- McGregor (2020) Sean McGregor. 2020. Preventing Repeated Real World AI Failures by Cataloging Incidents: The AI Incident Database. arXiv:2011.08512 [cs.CY] https://arxiv.org/abs/2011.08512
- Ng et al. (2022) DTK Ng, WY Luo, HMY Chan, and SKW Chu. 2022. An examination on primary students’ development in AI Literacy through digital story writing. Computers & Education: Artificial Intelligence 100054 (2022).
- Nielsen (1992) Jakob Nielsen. 1992. Finding usability problems through heuristic evaluation. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Monterey, California, USA) (CHI ’92). Association for Computing Machinery, New York, NY, USA, 373–380. https://doi.org/10.1145/142750.142834
- Nigam et al. (2024) Hitesh Nigam, Mayur Patwardhan, Lovekesh Vig, and Gautam Shroff. 2024. Acceleron: A Tool to Accelerate Research Ideation. arXiv preprint arXiv:2403.04382 (2024).
- Norman et al. (2021) Michael Norman, Vince Kellen, Shava Smallen, Brian DeMeulle, Shawn Strande, Ed Lazowska, Naomi Alterman, Rob Fatland, Sarah Stone, Amanda Tan, et al. 2021. CloudBank: Managed Services to Simplify Cloud Access for Computer Science Research and Education. In Practice and Experience in Advanced Research Computing. 1–4.
- Ntoa et al. (2021) Stavroula Ntoa, George Margetis, Margherita Antona, and Constantine Stephanidis. 2021. User experience evaluation in intelligent environments: A comprehensive framework. Technologies 9, 2 (2021), 41.
- OpenAI et al. (2024) OpenAI, Josh Achiam, Steven Adler, and et al. 2024. GPT-4 Technical Report. arXiv:2303.08774 [cs.CL] https://arxiv.org/abs/2303.08774
- Paas et al. (2008) FGWC Paas, Paul Ayres, and Mariya Pachman. 2008. Assessment of cognitive load in multimedia learning. Recent Innovations in Educational Technology That Facilitate Student Learning, Information Age Publishing Inc., Charlotte, NC (2008), 11–35.
- Paas (1992) Fred GWC Paas. 1992. Training strategies for attaining transfer of problem-solving skill in statistics: a cognitive-load approach. Journal of educational psychology 84, 4 (1992), 429.
- Pang et al. (2024) Rock Yuren Pang, Sebastin Santy, René Just, and Katharina Reinecke. 2024. BLIP: Facilitating the Exploration of Undesirable Consequences of Digital Technologies. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–18.
- Pecune et al. (2019) Florian Pecune, Shruti Murali, Vivian Tsai, Yoichi Matsuyama, and Justine Cassell. 2019. A model of social explanations for a conversational movie recommendation system. In Proceedings of the 7th International Conference on Human-Agent Interaction. 135–143.
- Pettersson et al. (2018) Ingrid Pettersson, Florian Lachner, Anna-Katharina Frison, Andreas Riener, and Andreas Butz. 2018. A Bermuda triangle? A Review of method application and triangulation in user experience evaluation. In Proceedings of the 2018 CHI conference on human factors in computing systems. 1–16.
- Reichheld (2003) Frederick F Reichheld. 2003. The one number you need to grow. Harvard business review 81, 12 (2003), 46–55.
- Rezwana and Maher (2023) Jeba Rezwana and Mary Lou Maher. 2023. User Perspectives on Ethical Challenges in Human-AI Co-Creativity: A Design Fiction Study. In Proceedings of the 15th Conference on Creativity and Cognition (Virtual Event, USA) (C&C ’23). Association for Computing Machinery, New York, NY, USA, 62–74. https://doi.org/10.1145/3591196.3593364
- Rivero and Conte (2017) Luis Rivero and Tayana Conte. 2017. A systematic mapping study on research contributions on UX evaluation technologies. In Proceedings of the XVI Brazilian symposium on human factors in computing systems. 1–10.
- Rodden et al. (2010) Kerry Rodden, Hilary Hutchinson, and Xin Fu. 2010. Measuring the user experience on a large scale: user-centered metrics for web applications. In Proceedings of the SIGCHI conference on human factors in computing systems. 2395–2398.
- Roselli et al. (2019) Drew Roselli, Jeanna Matthews, and Nisha Talagala. 2019. Managing Bias in AI. In Companion Proceedings of The 2019 World Wide Web Conference (San Francisco, USA) (WWW ’19). Association for Computing Machinery, New York, NY, USA, 539–544. https://doi.org/10.1145/3308560.3317590
- Sadowski et al. (2019) Caitlin Sadowski, Margaret-Anne Storey, and Robert Feldt. 2019. A software development productivity framework. Rethinking Productivity in Software Engineering (2019), 39–47.
- Schrepp et al. (2017) Martin Schrepp, Andreas Hinderks, and Jörg Thomaschewski. 2017. Design and evaluation of a short version of the user experience questionnaire (UEQ-S). International Journal of Interactive Multimedia and Artificial Intelligence, 4 (6), 103-108. (2017).
- Shankar et al. (2024) Shreya Shankar, JD Zamfirescu-Pereira, Björn Hartmann, Aditya G Parameswaran, and Ian Arawjo. 2024. Who Validates the Validators? Aligning LLM-Assisted Evaluation of LLM Outputs with Human Preferences. arXiv preprint arXiv:2404.12272 (2024).
- Shen et al. (2023) Hua Shen, Chieh-Yang Huang, Tongshuang Wu, and Ting-Hao Kenneth Huang. 2023. ConvXAI: Delivering Heterogeneous AI Explanations via Conversations to Support Human-AI Scientific Writing. In Companion Publication of the 2023 Conference on Computer Supported Cooperative Work and Social Computing (Minneapolis, MN, USA) (CSCW ’23 Companion). Association for Computing Machinery, New York, NY, USA, 384–387. https://doi.org/10.1145/3584931.3607492
- Shneiderman (2020) Ben Shneiderman. 2020. Bridging the gap between ethics and practice: guidelines for reliable, safe, and trustworthy human-centered AI systems. ACM Transactions on Interactive Intelligent Systems (TiiS) 10, 4 (2020), 1–31.
- Stankov et al. (2012) Lazar Stankov, Jihyun Lee, Wenshu Luo, and David J Hogan. 2012. Confidence: A better predictor of academic achievement than self-efficacy, self-concept and anxiety? Learning and individual differences 22, 6 (2012), 747–758.
- Stige et al. (2023) Åsne Stige, Efpraxia D Zamani, Patrick Mikalef, and Yuzhen Zhu. 2023. Artificial intelligence (AI) for user experience (UX) design: a systematic literature review and future research agenda. Information Technology & People (2023).
- Sun et al. (2024) Haowen Sun, Ruikun Zheng, Haibin Huang, Chongyang Ma, Hui Huang, and Ruizhen Hu. 2024. LGTM: Local-to-Global Text-Driven Human Motion Diffusion Model. In ACM SIGGRAPH 2024 Conference Papers. 1–9.
- Tamm et al. (2021) Yan-Martin Tamm, Rinchin Damdinov, and Alexey Vasilev. 2021. Quality metrics in recommender systems: Do we calculate metrics consistently?. In Proceedings of the 15th ACM Conference on Recommender Systems. 708–713.
- Tan et al. (2024) Felicia Fang-Yi Tan, Peisen Xu, Ashwin Ram, Wei Zhen Suen, Shengdong Zhao, Yun Huang, and Christophe Hurter. 2024. AudioXtend: Assisted Reality Visual Accompaniments for Audiobook Storytelling During Everyday Routine Tasks. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–22.
- Thomas et al. (2020) Paul Thomas, Daniel McDuff, Mary Czerwinski, and Nick Craswell. 2020. Expressions of style in information seeking conversation with an agent. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval. 1171–1180.
- Venable et al. (2012) John Venable, Jan Pries-Heje, and Richard Baskerville. 2012. A comprehensive framework for evaluation in design science research. In Design Science Research in Information Systems. Advances in Theory and Practice: 7th International Conference, DESRIST 2012, Las Vegas, NV, USA, May 14-15, 2012. Proceedings 7. Springer, 423–438.
- Verma et al. (2023) Shilpa Verma, Rajesh Bhatia, Sandeep Harit, and Sanjay Batish. 2023. Scholarly knowledge graphs through structuring scholarly communication: a review. Complex & intelligent systems 9, 1 (2023), 1059–1095.
- Vermeeren et al. (2010) Arnold POS Vermeeren, Effie Lai-Chong Law, Virpi Roto, Marianna Obrist, Jettie Hoonhout, and Kaisa Väänänen-Vainio-Mattila. 2010. User experience evaluation methods: current state and development needs. In Proceedings of the 6th Nordic conference on human-computer interaction: Extending boundaries. 521–530.
- Vinuesa et al. (2020) Ricardo Vinuesa, Hossein Azizpour, Iolanda Leite, Madeline Balaam, Virginia Dignum, Sami Domisch, Anna Felländer, Simone D Langhans, Max Tegmark, and Francesco Fuso Nerini. 2020. The role of artificial intelligence in achieving the Sustainable Development Goals. Nature communications 11, 1 (2020), 1–10.
- Wambsganss et al. (2020) Thiemo Wambsganss, Rainer Winkler, Matthias Söllner, and Jan Marco Leimeister. 2020. A conversational agent to improve response quality in course evaluations. In Extended Abstracts of the 2020 CHI conference on human factors in computing systems. 1–9.
- Wang et al. (2020) Dakuo Wang, Elizabeth Churchill, Pattie Maes, Xiangmin Fan, Ben Shneiderman, Yuanchun Shi, and Qianying Wang. 2020. From human-human collaboration to Human-AI collaboration: Designing AI systems that can work together with people. In Extended abstracts of the 2020 CHI conference on human factors in computing systems. 1–6.
- Wang et al. (2021) Di Wang, Ming Yin, and Gloria Mark. 2021. Do you trust me to trust you? Mental model transparency and personalized explainability in human-AI collaboration. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–13.
- Wang et al. (2018) Ruijie Wang, Yuchen Yan, Jialu Wang, Yuting Jia, Ye Zhang, Weinan Zhang, and Xinbing Wang. 2018. Acekg: A large-scale knowledge graph for academic data mining. In Proceedings of the 27th ACM international conference on information and knowledge management. 1487–1490.
- Wang et al. (2024) Zhan Wang, Lin-Ping Yuan, Liangwei Wang, Bingchuan Jiang, and Wei Zeng. 2024. Virtuwander: Enhancing multi-modal interaction for virtual tour guidance through large language models. In Proceedings of the CHI conference on human factors in computing systems. 1–20.
- Xu and Mulligan (2022) Horvitz Eric Zeng Eric Xu, Wesley and Deirdre K Mulligan. 2022. Demystifying AI among end users: A taxonomy of AI literacy. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. 1–19.
- Yang et al. (2020) Qian Yang, Aaron Steinfeld, Carolyn Rosé, and John Zimmerman. 2020. Re-examining whether, why, and how human-AI interaction is uniquely difficult to design. In Proceedings of the 2020 chi conference on human factors in computing systems. 1–13.
- Zhai et al. (2024) Chunpeng Zhai, Santoso Wibowo, and Lily D Li. 2024. The effects of over-reliance on AI dialogue systems on students’ cognitive abilities: a systematic review. Smart Learning Environments 11, 1 (2024), 28.
- Zheng and Huang (2023) Qingxiao Zheng and Yun Huang. 2023. ” Begin with the End in Mind”: Incorporating UX Evaluation Metrics into Design Materials of Participatory Design. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems. 1–7.
- Zheng et al. (2022) Qingxiao Zheng, Yiliu Tang, Yiren Liu, Weizi Liu, and Yun Huang. 2022. UX research on conversational human-AI interaction: A literature review of the ACM digital library. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. 1–24.
Appendix A Appendix
A.1. Definitions for Index
Paradigms:
Dyadic: One-on-one interaction between human and AI. Dyadic human-AI interaction, where end users interact with AI systems, and end users are not AI creators.
Polyadic: Multiple social interactions mediated by AI. Where AI sits between two or many types of end users, and these end users are not AI creators either.
Application domain: The specific field or sector in which the system is applied, such as education, healthcare, finance, urban planning, etc.
Modality: The form or channel through which users interact with the system, including Text-Based, Voice-Based, Multi-Modal (Text and Voice Based), Visual Based, Sensor Based, Haptic (touch) Based.
System Features: The characteristics and attributes of the system, including embodiment, level of intelligence, autonomy, adaptability, and learning capabilities.
Design Novelty: The uniqueness and innovativeness of the system’s design, indicating how novel the system is compared to existing technologies.
Design methods: Refers to the systematic approaches, strategies, and techniques employed in the creation and development of designs. These methods guide designers in the process of generating ideas, solving problems, and producing effective, functional, and aesthetically pleasing solutions. Example design methods can be:
Framework-based design (proof-of-concept designs guided by prior frameworks, models, or theories)
Participatory design
Need-finding interviews
Ideation workshops
User-Centered Design
Research through design
Design fiction
Human-Centered Design
Contextual design
Service design
Inclusive design
Ethnographic design
Critical design
Reflective design
Sustainable design
Wizard of Oz
Human-AI relationship types: The nature of the interaction and relationship between humans and the system, ranging from assistant, collaborator, tool, to advisor.
Stakeholders: Individuals or groups who have an interest in, are affected by, or can influence the design, development, deployment, and use of the system. This includes:
Primary Users: Main direct users.
Secondary Users: Occasional or indirect users.
Tertiary Users: Affected non-users.
Designers and Developers: System creators and maintainers.
Organizations: Entities managing or funding the system.
Policymakers: Regulators setting guidelines.
Support Teams: Maintenance personnel.
Researchers: Those studying and evaluating the system.
Social scale: The scale at which the system operates, such as individual, group, organizational, societal, etc.
Theoretical Frameworks: The underlying theories and concepts that support the development and application of the system, which may include theories from cognitive science, social science, computer science, etc.
| Node Type | Attributes | Explanation |
| Paper |
Title
Narrative Indexes CommunityID Narrative Embedding Index Embedding |
The research paper title
The narrative/abstract of the research paper The results corresponding to each index category The communityID to which the research paper is classified into The embedding of the research paper narrative The embedding of each index value |
| Metric |
Metric Name
Metric Type |
The metric name
The category to which the metric is assigned |
| Outcome |
Outcome Achieved
Paper Title Paper Citation Metric Method Metric Usage Metric Challenges |
Metric outcome/result
The research paper title The reason for using the metric Method used to evaluate the metric How the metric is used in the research paper The challenges or research gap identified in the research paper |
A.2. User Study Survey
Background collection
What’s your name?
What’s your email address?
What’s your age? (optional)
How do you identify your gender? Male, Female, Prefer not to say, Other, please specify
What degree program are you currently enrolled in? Bachelor’s Degree, Master’s Degree, Doctorate (PhD or equivalent), Other, please specify
UX research experience: How many years of UX research experience do you have? None, less than 1 year, 2-3 years, 3-4 years, 5+ years
Coding experience: How many years of coding experience do you have? None, less than 1 year, 2-3 years, 3-4 years, 5+ years
AI literacy (Ng et al., 2022): How strongly do you agree or disagree with the following statements?
I can operate AI applications in everyday life.
I can use AI meaningfully to achieve my everyday goals.
I can assess what the limitations and opportunities of using an AI are.
AI Assistance in UX Research: Currently, what types of research tasks do you use AI (e.g., ChatGPT) to assist with? Research Ideation, Literature review, Evaluation plan development, Data analysis, Prototyping and design, Usability testing, Coding, Writing, Other please specify
Survey 1 and 2 Initial/Final UX Study Plan: We are going to ask about your perceptions of the proposed UX plans for your project.
How do you feel about the following statements about your proposed UX plans? Strongly disagree, Disagree, Neither agree nor disagree, Agree, Strongly agree
Novelty (Dean et al., 2006): My proposed UX plans are novel.
Relevance (Dean et al., 2006): My proposed UX plans are relevant to my research goal.
Efficiency (Venable et al., 2012): My proposed UX plans are efficient in supporting my research goals.
Completeness (Dean et al., 2006): My proposed UX plans are complete.
Clarity (Dean et al., 2006): My proposed UX plans are clear.
Concreteness (Dean et al., 2006): My proposed UX plans are concrete.
Quality (Dean et al., 2006): My proposed UX plans are of high quality.
Feasibility (Dean et al., 2006): My proposed UX plans is workable and can be easily implemented.
Rigorous (Venable et al., 2012): My proposed UX plans are methodologically rigorous.
Risk considerations (Venable et al., 2012): My proposed UX plans are in consider of potential issues.
Confidence (Stankov et al., 2012): My proposed UX plans are making me feel confident about them.
Survey 3 System Features: We are going to ask about your perceptions of each feature in our system. How do you feel about the following statements regarding the Index/ Metric Graph/ Metric List/ Outcome/ Risk feature? (We repeated the same set of questions for each feature.) Strongly disagree, Disagree, Neither agree nor disagree, Agree, Strongly agree
I enjoy using this feature. (Davidson et al., 2023)
I will use this feature in the future if it is publicly available (Everard and Galletta, 2005).
I find this feature useful (Davis, 1989).
I trusted the suggestions and decisions provided by it (Everard and Galletta, 2005).
It took significant mental effort to understand this feature (Paas, 1992).
I was more interested in exploring and developing my research projects/UX evaluation plan/UX outcomes after using it (Chu et al., 2010).
This feature contributes to revising my UX evaluation plan.
This feature contributes to revising my UX outcomes.
A.3. Data Analysis
| ID | Education Level | UX Research Experience | Coding Experience | AI Literacy Measured by (Ng et al., 2022) | AI Usage in Research |
| P1 | Master’s | 3-4 years | 1-2 years | 4.33 | Research Ideation, Evaluation Plan Development, Coding, Writing, etc. |
| P2 | Master’s | 3-4 years | 5+ years | 5.00 | Research Ideation, Evaluation Plan Development, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P3 | Doctorate | 1-2 years | 3-4 years | 4.67 | Literature Review, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P4 | Master’s | 3-4 years | 5+ years | 4.67 | Research Ideation, Literature Review, Data Analysis, Coding, Writing, etc. |
| P5 | Postdoc | 1-2 years | 3-4 years | 4.00 | Literature Review, Usability Testing, Coding, Writing, etc. |
| P6 | Doctorate | 1-2 years | 3-4 years | 5.00 | Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P7 | Doctorate | 1-2 years | 5+ years | 4.00 | Data Analysis, Coding, Writing, etc. |
| P8 | Doctorate | 3-4 years | 5+ years | 5.00 | Literature Review, Data Analysis, Coding, Writing, etc. |
| P9 | Doctorate | 1-2 years | 3-4 years | 4.00 | Research Ideation, Coding, Writing, etc. |
| P10 | Doctorate | 3-4 years | 3-4 years | 4.00 | Data Analysis, Coding, Writing, etc. |
| P11 | Doctorate | 3-4 years | 5+ years | 3.67 | Research Ideation, Data Analysis, Coding, Writing, etc., Other, please specify |
| P12 | Master’s | 1-2 years | 3-4 years | 5.00 | Research Ideation, Literature Review, Evaluation Plan Development, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P13 | Doctorate | 1-2 years | 3-4 years | 4.00 | Coding, Writing, etc. |
| P14 | Bachelor’s | 3-4 years | 3-4 years | 5.00 | Literature Review, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P15 | Doctorate | 3-4 years | 5+ years | 4.00 | Coding, Writing, etc. |
| P16 | Doctorate | Less than 1 year | 5+ years | 4.33 | Other, debugging front-end UI development |
| P17 | Doctorate | 3-4 years | 5+ years | 4.67 | Other, integrating various LLMs into our embodied robots to support dynamic interactions w/ users |
| P18 | Doctorate | 3-4 years | 5+ years | 5.00 | Research Ideation, Literature Review, Prototyping and Design, Data Analysis, Coding, Writing, etc. |
| P19 | Master’s | 3-4 years | 5+ years | 4.00 | Prototyping and Design, Coding, Writing, etc. |
| Constructs | Before Median | Before IQR | After Median | After IQR | Test Statistic | Effect Size |
| Novelty | 3 | 2 | 4 | 1 | 22 | 0.51 |
| Relevance | 5 | 1 | 5 | 0.5 | 8 | 0.56 |
| Efficiency∗ | 4 | 0 | 5 | 1 | 12 | 0.74 |
| Completeness∗∗ | 3 | 2 | 4 | 1 | 16 | 0.79 |
| Clarity∗∗ | 4 | 1 | 4 | 1 | 4 | 0.90 |
| Concreteness∗ | 4 | 1 | 4 | 1.5 | 24 | 0.74 |
| Quality∗∗ | 3 | 1 | 4 | 0.5 | 11 | 0.82 |
| Workability∗ | 4 | 2 | 5 | 2 | 16.5 | 0.64 |
| Rigor∗ | 3 | 2 | 4 | 1.5 | 15 | 0.75 |
| Ethics∗∗ | 3 | 2.5 | 4 | 1.5 | 10 | 0.81 |
| Trust∗ | 4 | 1 | 4 | 1 | 5 | 0.82 |
| Confidence∗ | 3 | 1 | 4 | 1 | 18 | 0.70 |